i am stuck on these problems...can someone help?
1) (sqrt of x) - 1 = 3
2) (sqrt of x^3) = 8

Answer by Nate(3500)

For x^(a/b)
the numerator (a) determines if the number is squared or cubed
the denominator (b) determines the type of root (ex. cubed root)

1) x^(1/2) - 1 = 3 original
x^(1/2) = 4 add the one
x = 16 square the whole equation to get x^1

2) x^(3/2) = 8
x = 4 multiply the exponents by 3/2 to get x^1